### v1

# UrbanKGent: A Unified Large Language Model Agent Framework for Urban Knowledge Graph Construction

Abstract.
Urban knowledge graph has recently worked as an emerging building block to distill critical knowledge from multi-sourced urban data for diverse urban application scenarios. Despite its promising benefits, urban knowledge graph construction (UrbanKGC) still heavily relies on manual effort, hindering its potential advancement. This paper presents UrbanKGent, a unified large language model agent framework, for urban knowledge graph construction. Specifically, we first construct the knowledgeable instruction set for UrbanKGC tasks (such as relational triplet extraction and knowledge graph completion) via heterogeneity-aware and geospatial-infused instruction generation. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectories distilled from GPT-4. Through hybrid instruction fine-tuning with augmented trajectories on Llama-2-13B, we obtain the UrbanKGC agent, UrbanKGent-13B. We perform a comprehensive evaluation on two real-world datasets using both human and GPT-4 self-evaluation. The experimental results demonstrate that UrbanKGent-13B not only can significantly outperform 21 baselines in UrbanKGC tasks, but also surpass the state-of-the-art LLM, GPT-4, by more than 10% with approximately 20 times lower cost. We deploy UrbanKGent-13B to provide online services, which can construct an UrbanKG with thousands of times richer relationships using only one-fifth of the data compared with the existing benchmark. Our data, code, and opensource UrbanKGC agent are available at https://github.com/usail-hkust/UrbanKGent. Refer to caption Figure 1. The framework overview of UrbanKGent.
urban knowledge graph, knowledge graph construction, large language model agent
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligence; Natural language processing; Information extraction
1. Introduction
Urban Knowledge Graph (UrbanKG) aims to model intricate relationships and semantics within a city by extracting and organizing urban entities (e.g., POIs, road networks, etc.) into a multi-relational heterogeneous graph. As an emerging building block, multi-sourced urban data are widely used to construct an UrbanKG to provide critical knowledge for various knowledge-enhanced urban downstream tasks, such as traffic management, pollution monitoring, and emergency response (Sun et al., 2017; Chen et al., 2016; Cheng et al., 2018). UrbanKG has gradually become an essential tool of the modern smart city.
In prior literature, many efforts have been devoted to urban knowledge graph construction (UrbanKGC) using massive urban data sources. In particular, one commonly used approach (Wang et al., 2021; Liu et al., 2022; Ning et al., 2023) is to extract entities from structured urban data (e.g., geographic data, city sensor data, and traffic data) and define the relationships between obtained urban entities based on manually designed rules or patterns. However, these approaches suffer heavy reliance on a deep understanding of the application domain and are labor-intensive. Recently, inspired by the success of the Large Language Models (LLMs) in knowledge graph construction (Xie et al., 2023; Zheng et al., 2021; Yao et al., 2023), the LLMs have been adopted to improve UrbanKGC. For instance, GeoLM (Wang et al., 2023b) is proposed for spatial grounded entity recognition and relation extraction. K2 (Deng et al., 2023) retrains the Llama-2-7B model on a huge annotated geoscience text corpus for geospatial relation extraction. Nevertheless, these works rely on annotated corpus and model retraining, which may discourage researchers from adopting it for their own work and thus limit the application of UrbanKG.
LLM agent (et al, 2023a, b) has recently emerged and shown remarkable zero-shot capability for autonomous domain-specific task completion. For example, Voyager (Wang et al., 2023c) is a LLM-powered agent for zero-shot game exploration without re-training, and LLMLight (Gur et al., 2023) is a traffic signal control agent with zero-shot LLM reasoning ability. These studies motivate us to construct tailored LLM agents to address the aforementioned limitations in UrbanKG construction.
Figure 2. Illustrative example of urban relational triplet extraction and knowledge graph completion. (a) The heterogeneous relationship understanding limitation of LLMs can be addressed by injecting prior urban knowledge into instruction. (b) The geospatial computing limitation of LLMs can be alleviated by invoking external geospatial tools.
In fact, constructing an LLM agent compatible with various UrbanKGC tasks is a non-trivial problem due to the following two challenges: (1) Challenges 1: How to adapt LLMs for UrbanKGC? LLMs may not align well with the specific task due to the gap (Fu et al., 2023b) between the natural language processing corpus for training LLMs and the domain-specific corpus in urban domain (Lin et al., 2023). For example, the urban text data is usually heterogeneous and contains multifaceted urban knowledge (e.g., spatial, temporal, and functional aspects) (Deng et al., 2023). As shown in Figure 2(a), the text description of ”Columbia University” reflects its geographic spatial locations (i.e., spatial relationship), construction timelines (i.e., temporal relationship), and how it provides educational service for the city (i.e., functional relationship). LLMs may require a tailored alignment to understand heterogeneous urban relationships to extract these urban spatial, temporal, and functional relations accurately. (2) Challenges 2: How to improve the capacity of LLMs for UrbanKGC? The effectiveness of LLMs for urban knowledge graph construction is restricted by their feeble numerical computation capacity (Frieder et al., 2023; Zhang et al., 2023b), leading to their disability in complex geospatial relationship extraction (Bhandari et al., 2023; Mooney et al., 2023). However, the urban geospatial relationship plays a vital role in urban semantic modeling (Li et al., 2023c) and has been widely incorporated in previous UrbanKGs (Ning et al., 2023; Liu et al., 2023). As can be seen in Figure 2, extracting ”disconnected” relation between the geo-entity ”Columbia University” and ”Empire State Building” is useful for urban geo-semantic modeling. Accurately extracting such geospatial relationship demands necessary geospatial computing (e.g., utilizing latitude and longitude for distance calculation) and reasoning (i.e., deriving calculation results for geospatial relation reasoning) capabilities. It is appealing to improve the geospatial computing and reasoning ability (e.g., invoking external tools for calculation) of LLMs to satisfy the UrbanKGC task requirement.
To address the aforementioned challenges, in this study, we propose UrbanKGent, a unified LLM agent framework for automatic UrbanKG construction. Figure 1 illustrates the overview of UrbanKGent. For a given city, we first generate a knowledgeable instruction set for UrbanKGC tasks (e.g., relational triplet extraction and knowledge graph completion) from urban geographic and text data sources. By heterogeneity-aware and geospatial-infused instruction generation, as shown in Figure 2(a), various urban spatiotemporal relationship knowledge can be encoded into the instruction, which facilitates alignment between LLMs with the target UrbanKGC tasks. Moreover, we propose a tool-augmented iterative trajectory refinement module to enhance and refine the trajectory derived by distilling GPT-4 with the above constructed instructions. Based on geospatial tool augmentation and self-refinement, the deficiency of LLMs in geospatial computing and reasoning could be alleviated, and unfaithful trajectories could be filtered out. Finally, we perform hybrid instruction fine-tuning based on the enhanced and refined trajectories on Llama 2-7B and Llama 2-13B variants (Touvron et al., 2023) by using LoRA (Hu et al., 2021). The obtained agent, UrbanKGent-13B, is feasible for completing multiple UrbanKGC tasks in a cost-effective way (i.e., no extra GPT-API cost needed).
We conduct comprehensive experiments on two UrbanKGC tasks in two metropolises (New York City and Chicago) using both human evaluation and GPT-4-based self-evaluation. The empirical results validate the effectiveness of the proposed LLM agent for completing various UrbanKGC tasks. Moreover, we deploy the constructed agent, UrbanKGent-13B, online for UrbanKGC service, which can extract the same scale of triplets and entities of existing UrbanKG benchmark (Ning et al., 2023) using only one-fifth of data, and even expand the types of relations by thousands of times.
Our contributions are summarized as follows:
  • We propose the first UrbanKGC agent construction framework UrbanKGent, and release the UrbanKGent-13B to provide real-world UrbanKGC service [^1] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html, which offers new opportunities to advance UrbanKG studies.
  • We propose a knowledgeable instruction generation module and a tool-augmented iterative trajectory refinement method, which align LLMs to UrbanKGC tasks and compensate for their geospatial computing and reasoning inability.
  • Extensive experiments on two real-world datasets validate the effectiveness of the proposed framework and uncover its exceptional performance across various UrbanKGC tasks.
2. UrbanKGC Data Description
This section introduces the process of data collection and preprocessing details.
2.1. Data Collection
We first acquire urban knowledge for two large cities New York City and Chicago from two data sources. Table 1 summarizes the statistics of the raw datasets.
2.1.1. Geographic Data
The geographic data provides critical urban spatial structure information and functional semantics, which has been widely used in previous UrbanKG studies (Ning et al., 2023; Liu et al., 2023, 2021; Zhang et al., 2023a).
Area-Of-Interst (AOI) Data. AOI data describes the urban spatial area structure, including urban commercial areas (e.g., shopping centers), residential areas (e.g., communities and neighbors), and so on. In this work, we first follow UUKG (Ning et al., 2023) to acquire the AOI name and geometry value from NYC Gov [^2] https://www.nyc.gov/ and CHI Gov [^3] https://www.chicago.gov/. Next, we use the AOI name to search for their text description from Wikipedia website [^4] https://www.wikipedia.org/. If there is no direct match, we then retrieve the Wikipedia pages that mention the AOI name and collect the related paragraphs. Each AOI record contains an AOI name, a polygon geometry value, and a text description. For example, [”Jamaica Bay”, polygon (-73.86 40.58, …), ”Jamaica Bay is an estuary …”] is the record of the AOI ”Jamaica Bay” with geometry value and text description.
Road Network Data. Road data describes the urban spatial transportation network, including urban motorways, overpasses, and so on. We first follow (Ning et al., 2023) to obtain the road name and geometry value from Open Street Map (OSM) [^5] https://www.openstreetmap.org/. Then, following the same text acquisition operation in AOI data, we crawl the textual description of each road record from Wikipedia. Each road record contains a road name, a linestring geometry value, a road type and a text description. For example, [”Central Park Avenue”, linestring (-73.87 40.90, …), primary, ”Central Park Avenue is a boulevard in …”] describes the primary road named ”Central Park Avenue” with a linestring geometry value and its textual description.
Point-Of-Interest (POI) Data. POI data represents different urban functions (e.g., residential and commercial), which have been widely adopted in many recent UrbanKG works (Liu et al., 2023; Wang et al., 2021; Ning et al., 2023). We first follow (Ning et al., 2023) to obtain the POI name, and geometry value from OSM. Then the textual description of each POI record could be crawled from Wikipedia following the similar process. Each POI record contains a POI name, a coordinate geometry, a POI type, and a text description. For example, [”Trump World Tower”, coordinate (-73.96 40.75), residential, ”Trump World Tower is a residential condominium …”] is the record of the POI ”Trump World Tower”.
2.1.2. Text Data
The text data provides rich contextual knowledge of the city space from different perspectives (e.g., the spatial context) (Deng et al., 2023), and it plays an important role in geospatial understanding. In this work, we collect two types of text corpus.
Review Data. The review of urban places provides useful commercial information that citizens use to make business decisions (Cho et al., 2011), which plays a critical role in urban knowledge distillation. We collect review data from Google Map [^6] https://www.google.com/maps. Specifically, we first manually split the city into multiple rectangular regions, then we utilize the Google Map API to query the places contained within each region and their reviews. Each review record contains a place name, a coordinate geometry value, a rating, and a text review. For example, [”Lifestyles Academy Inc”, coordinate (-87.87 41.65), 4.9, ”Very nice organization and …”] is the review record of place ”Lifestyles Academy Inc”.
Web Page Data. The web page data works as the general text corpus for the city, and it contains rich geoscience knowledge that has been utilized in recent urban entity and relation extraction studies (Deng et al., 2023). We collect web page data from the Google search engine. Specifically, we first input the name of the crawled AOI, Road, and POI record into Google. Then we concatenate the textual sentences of the top 10 retrieved web pages. Each web page record contains a long urban text description.
Table 1. The statistics of raw datasets.
Dataset Description New York City Chicago
Geographic Data # of AOI 183 80
# of road 6,397 1,893
# of POI 5,369 5,658
Text Data # of review 16,360 13,627
# of web page 11,596 7,283
2.2. Data Preprocessing
Before constructing the UrbanKGC dataset, we first preprocess the raw datasets. Specifically, we filter out AOIs, roads, POIs, reviews, and web pages whose crawled textual descriptions are null value, too short (e.g., less than ten word description) or meaningless (e.g., just repeating the POI name). In addition, we remove irrelevant information from the text description, such as non-English characters, non-ASCII gibberish, website addresses, and so on. More details can be found in Appendix A.
3. Preliminary
This section presents the UrbanKGC task definition and provides task analysis.
3.1. Task Definition and Problem Formulation
Before diving into the technical details, we first introduce the definition of urban knowledge graph (UrbanKG):
Definition 1 ().
UrbanKG. The UrbanKG is defined as a multi-relational graph $\mathcal{G}$ = ($\mathcal{E}$, $\mathcal{R}$, $\mathcal{F}$), where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{F}$ is the set of urban entities, relations and facts, respectively. In particular, facts are defined as =$\mathcal{F}=\{\langle h,r,t\rangle\mid h,t\in\mathcal{E},r\in\mathcal{R}\}$, where each triplet ${\langle h,r,t\rangle}$ describes head entity ${h}$ is connected with tail entity ${t}$ via relation ${r}$.
The UrbanKG encodes diverse urban semantic knowledge by connecting urban entities into a multi-relational graph. This work aims to construct an UrbanKG from collected unstructured text data. We decompose the UrbanKG construction (UrbanKGC) process into two sequential knowledge graph construction tasks, namely relational triple extraction (Zheng et al., 2021) and knowledge graph completion (Yao et al., 2023). We first provide the basic definition for these two subtasks, and then introduce the problem formulation of this work.
3.1.1. Task Definition
We give the basic task definition as follow.
Relational Triplet Extraction (RTE). Given the unstructured texts, this task achieves joint extraction of entities and their relations (Zheng et al., 2021) which are in the form of a triplet ${\langle h,r,t\rangle}$. For instance, given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, this task aims to identify two entities ”Columbia University” and ”New York City” and their relation ”locate-in”, described as triplet ¡Columbia University, locate-in, New York City¿.
Knowledge Graph Completion (KGC). Given a head entity $h$ and a tail entity $t$, this task is to predict the missing relation between them (Yao et al., 2023). For instance, given the head entity ”Columbia University” and the tail entity ”Empire State Building”, this task is to predict that their missing relation, e.g., ”disconnected”, described as triplet ¡Columbia University, disconnected, Empire State Building¿.
3.1.2. Problem Formulation
Given the urban unstructured text data, the desired output is an UrbanKG $\mathcal{G}$. In this paper, this problem is decomposed into two sequential subtasks: (1) Relational Triplet Extraction: the first task extracts relational triplet ${\langle h,r,t\rangle}$ from the urban text data. The output of RTE task is $\mathcal{G}_{1}$ $\mathcal{E}$, $\mathcal{R}_{1}$, $\mathcal{F}_{1}$), where $\mathcal{E}$ and $\mathcal{R}_{1}$is the set of extracted entities and relations, while $\mathcal{F}_{1}$ is the set of extracted triplets. (2) Knowledge Graph Completion: for the given head entity $h$ and tail entity $t$ in $\mathcal{G}_{1}$, the second task is to predict the geospatial relationship [^7] We follow GeoLM (Li et al., 2023c) to provide five RCC relationship (Mani et al., 2010) candidates: Disconnection (DC), external connection (EC), equality (EQ), partial overlap (PO), and tangential and non-tangential proper parts (IN). Details are in Appendix A. between them. The output of this task is $\mathcal{G}_{2}$ $\mathcal{E}$, $\mathcal{R}_{2}$, $\mathcal{F}_{2}$), where $\mathcal{R}_{2}$ and $\mathcal{F}_{2}$ is the set of completed relations and triplets. By sequentially completing the above two tasks, we can obtain the constructed UrbanKG $\mathcal{G}$ = ($\mathcal{E}$, $\mathcal{R}_{1}\cup\mathcal{R}_{2}$, $\mathcal{F}_{1}\cup\mathcal{F}_{2}$).
3.2. Quantitative Task Analysis
Figure 3. Quantitative performance analysis of prompting GPT-4 for RTE and KGC tasks. The result is obtained by comparing 50 GPT-4’s outputs with the human’s annotation.
As shown in Figure 2, we qualitatively find that LLMs lack urban heterogeneous relationship understanding ability and experience in geospatial computing and reasoning difficulty when adopting it for UrbanKGC tasks. This subsection presents a quantitative analysis of these two challenges.
Heterogenous Relationship Understanding. The ability to understand heterogeneous relationships is ubiquitous in distilling knowledge from the massive urban corpus. For example, the text description in Figure 2 illustrates a place from spatial location, temporal time, and functional aspects. Capturing these heterogeneous semantics is important for urban knowledge distillation. We perform quantitative analysis by randomly sampling 50 urban text data and then prompt GPT-4 to complete relational triplet extraction by providing only the basic task description. As shown in Figure 3(a), we find that the LLMs experience serious misjudgment (i.e., extract wrong triplets or miss the triplet) on urban spatial, temporal, and functional triplet extraction. This indicates the limited capacity of LLMs to understand heterogeneous relationships.
Geospatial Computing and Reasoning. Geospatial computing and reasoning techniques are widely used in many previous UrbanKG studies (Ning et al., 2023; Liu et al., 2023) for urban geospatial relation extraction. In recent works (Mooney et al., 2023; Manvi et al., 2023), the geospatial skills of LLMs have also been demonstrated to lack geospatial awareness and reasoning ability (Bhandari et al., 2023). To identify potential limitations, we quantitatively investigate how LLMs can perform on geospatial relation completion tasks. Specifically, we construct 100 head and tail entity pairs, covering five geospatial relations in the KGC task, and then prompt GPT-4 to predict with basic task description and geospatial relation candidates. As shown in Figure 3(b), we find that GPT-4 performs poorly on five geospatial relation completion. This further validates the disability of LLMs in geospatial computing and reasoning.
4. UrbanKGC Agent Construction
This section presents the proposed UrbanKGC agent construction framework.
4.1. Overview
The overall pipeline of the UrbanKGent framework is illustrated in Figure 4, which consists of three major components: (1) Knowledgeable Instruction Generation consists of the heterogeneity-aware and geospatial-infused instruction generation modules for aligning LLMs to UrbanKGC tasks. (2) Tool-augmented Iterative Trajectory Refinement proposes geospatial tool interface invocation and iterative self-refinement mechanisms to enhance and refine generated trajectory. (3) Hybrid Instruction Fine-tuning fine-tune LLMs based on the refined trajectories for cost-effectively completing diverse UrbanKGC tasks.
4.2. Knowledgeable Instruction Generation
We first construct the knowledgeable instruction to adopt LLMs for two UrbanKGC tasks, including relational triplet extraction (RTE) and knowledge graph completion (KGC). Figure 4(a) illustrates the overview of the instruction construction process of these two tasks.
4.2.1. Heterogeneity-aware Instruction Generation for Relational Triplet Extraction.
As discussed in Section 3, the urban text contains diverse heterogeneous relationships, thus we consider multiple views with both urban entity and relation definition for relational triplet extraction. In particular, we construct a multi-view instruction template for the urban relational triplet extraction, including spatial view, temporal view, and functional view. Each view is a multi-turn question-answer dialog (Wei et al., 2023) consisting of entity recognition, relation extraction, and triplet extraction module.
For the spatial view, we devise a two-turn dialog to align LLMs for spatial triplet extraction. In the first turn, we inject spatial entity and relation definition into the instruction template to guide LLMs to understand spatial characteristics and then extract potential spatial entities (e.g., University) and relations types (e.g., locate-in). In the second turn, the extracted types are explicitly fed into the instruction template for spatial triplet extraction. Intuitively, the spatial view allocates dedicated urban knowledge for LLMs to extract urban spatial relationships. Similarly, we construct the temporal view and functional view for corresponding temporal and functional triplet extraction, independently. We provide the detailed instruction template in the Appendix B.
4.2.2. Geospatial-infused Instruction Generation for Knowledge Graph Completion.
Despite heterogeneity-aware instruction enabling LLMs to extract urban triplets from various perspectives, the geospatial relationship between geospatial entities cannot be directly extracted. Therefore, we introduce a geospatial-infused instruction generation module to guide LLMs to complete missing geospatial relationships.
First, we incorporate the geometry information (i.e., the latitude and longitude) of geo-entities into instruction, so that the LLMs can utilize these geospatial values for relation inference. Second, we add the geospatial relationship definition to the instruction to guide LLMs in understanding the geospatial relationship definition. Intuitively, LLMs can refer to this geospatial knowledge and make practical solutions for the knowledge graph completion task. We provide the detailed instruction template in the Appendix B.
4.3. Tool-augmented Iterative Trajectory Refinement
Then, we present the trajectory generation to further augment and refine the constructed UrbanKGC instructions.
4.3.1. Trajectory Generation
With the initial UrbanKGC instructions constructed, the following step is to generate reasoning trajectories (Zeng et al., 2023), which will be used to fine-tune LLMs tailored to UrbanKGC task. Specifically, we follow FireAct (Chen et al., 2023) and use Chain-of-Thought (CoT) (Wei et al., 2022), a gradient-free technique, to prompt GPT-4 (i.e., add prompt trigger ”Let’s think step by step” at the end of RTE and KGC instructions template) to generate the reasoning trajectories for UrbanKGC tasks.
The generated CoT trajectories could provide a step-by-step reasoning solution for UrbanKGC tasks. Nevertheless, the complex geospatial relationships cannot be easily extracted as discussed in Section 3 and recent geospatial reasoning works (Manvi et al., 2023; Mooney et al., 2023; Bhandari et al., 2023). Therefore, we introduce a tool invocation module to guide LLMs to invoke tailored external geospatial tools (Cai et al., 2023) to enhance their geospatial computing and reasoning capacity for UrbanKGC tasks.
Figure 4. An overview of UrbanKGent Construction.
4.3.2. Tool Invocation for Trajectory Augmentation
We conduct two sequential procedures: tool invocation for geospatial computing support and trajectory deliberation for reasoning enhancement.
Tool Invocation. First, we construct a geospatial reasoning toolkit (e.g., distance calculation, eight interfaces in total shown in Table 8) by prompting GPT-4 for self-programming. Then, we construct tailored prompts to guide LLMs to invoke these interfaces. Specifically, the prompt is concatenated with an illustrative description of the function of each geospatial tool and a task instruction (i.e., ”Which types of tool interface you need”). Intuitively, the external tool allocates necessary calculation results for LLMs to infer missing geospatial relation. The detailed toolkit description can be found in Appendix C.
Trajectory Deliberation. After manipulation with external tools, we prompt LLMs to refine uncertain reasoning steps based on these obtained manipulation results. Specifically, we construct the prompt by concatenating with manipulation results (e.g., the distance and geohash value of geo-entity) and a task instruction (i.e., ”Please refine your reasoning process”). After feeding the prompt into GPT-4, the enhanced trajectory is obtained. Detailed prompt information can be found in Appendix C.
4.3.3. Iterative Trajectory Self-refinement
Despite tool-augmented deliberation improving the geospatial computing and reasoning ability of LLMs, enhanced trajectories may not all be faithful (Hong et al., 2023). To alleviate potential error and ensure the trajectory quality (Huang et al., 2023), we refine these trajectories via an iterative self-refinement mechanism (Madaan et al., 2023). Specifically, we iterate two sequential blocks: (i) Trajectory verifier: given the trajectory, the verifier aims to provide feedback for refining the reasoning process; (ii) Trajectory updater: given the trajectory and feedback, the updater will further refine the current trajectory based on the feedback.
Trajectory Verifier. We construct a tailored prompt to ask LLMs to generate feedback. Specifically, we use a simple but effective trigger (”Judge whether all extracted triplets are correct and provide improvement suggestion”) to prompt LLMs to provide feedback. If the trajectory no longer requires modification, we let LLMs respond with ”This is a faithful trajectory”. Such a verification step lets LLMs make reflections and improve the correctness of the trajectory.
Trajectory Updater. The updater utilizes provided feedback to refine the current trajectory via prompt trigger ”Follow suggestion to refine the reasoning process”. Intuitively, the feedback may address multiple aspects (e.g., missed triplet in the RTE task or unfaithful reasoning process in the KGC task) of the unfaithful trajectories.
We iterate the trajectory verifier and updater until the predefined stopping condition is satisfied. The stopping condition is determined by either meeting the maximum number of iterations (we set it at three to avoid excessive cost) or when the verifier confirms all trajectories are faithful. Upon meeting the stopping condition, we use the last refined trajectory for further fine-tuning. Detailed prompt information can be found in Appendix C.
4.4. Hybrid Instruction Fine-Tuning
To construct a cost-effective UrbanKGC agent, we further utilize trajectories (generated by GPT-4) to fine-tune a smaller open-source LLM for faster inference speed and lower cost (i.e., prompting GPT-4 for UrbanKGC is expensive). Specifically, we finetune the LLM via the mixed-task instruction-tuning strategy (Zeng et al., 2023). The goal is to enhance the LLMs’ capabilities in diverse UrbanKGC tasks.
4.4.1. Mixture Training
Set the base language model as $\mathcal{M}$, and $\mathcal{P}_{\mathcal{M}}\left(y\mid x\right)$ represents the probability distribution of response $y$ given instruction $x$. We consider the trajectory set on two UrbanKGC tasks, i.e., $\mathcal{D}_{RTE}$ and $\mathcal{D}_{KGC}$. Since both the instruction and the target output are formatted in natural language, we can unify the training into an end-to-end sequence-to-sequence way. Formally, the optimization process aims to minimize the loss of language model $\mathcal{M}$ as follows:
(1) $\mathcal{L}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{RTE}}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right]+\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{KGC }}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right],$
where $x$ and $y$ represent the instruction input and instruction output in the trajectory, respectively.
4.4.2. Training Setup
We choose the chat version of open-sourced Llama 2-7/13B (Touvron et al., 2023) as our backbone models, and fine-tune Llama using Low-Rank Adaptation (LoRA) strategy (Hu et al., 2021). Through LoRA, the training process could be done by tuning only one-thousandth of full LLM parameters.
4.5. Inference on UrbanKGC Task
Via hybrid instruction fine-tuning, the LLM can be trained to follow the instructions to finish the UrbanKGC task. After obtaining the UrbanKGent, we can prompt it to complete UrbanKGC tasks by following the pipeline shown in Figure 4.
For the relational triplet extraction task, we sequentially execute entity recognition, relation extraction, and relational triplet instruction generation. We then perform iterative self-refinement and finally output the extracted triplets. For the knowledge graph completion task, we sequentially execute KGC instruction generation, external tool augmentation, iterative self-refinement block, and finally output the completed triplets.
Table 2. The statistics of constructed UrbanKGC dataset.
Dataset NYC-Instruct NYC CHI-Instruct CHI
Records RTE 232 2,089 122 1,102
KGC 232 2,080 122 1,101
5. Experiments
We evaluate the proposed method on two real-world UrbanKGC datasets, and aim to answer the following research questions:
  • RQ1: How does the constructed UrbanKGent perform compared with existing baselines and LLM prompting paradigms on UrbanKGC tasks?
  • RQ2: How do different components (e.g., the knowledgeable instruction generation) affect the performance?
  • RQ3: How can the constructed UrbanKGent provide application service to real-world scenarios?
Table 3. The main results of relational triplet extraction (RTE) and knowledge graph completion (KGC). We report the accuracy (acc) and confidence for GPT evaluation on two datasets, and report accuracy (acc) for the Human evaluation approach. The best baseline performance is underlined.
Type Models NYC CHI
GPT (acc/confidence) Human (acc) GPT (acc/confidence) Human (acc)
RTE KGC RTE KGC RTE KGC RTE KGC
Pretrained Language Models KG-BERT - 0.24/3.15 - 0.23 - 0.19/4.12 - 0.24
KG-T5 - 0.21/4.02 - 0.21 - 0.15/3.98 - 0.24
RelationPrompt 0.12/3.38 - 0.12 - 0.21/3.53 - 0.18 -
Zero-shot Reasoning Llama-2-7B 0.14/1.98 0.18/3.75 0.16 0.18 0.26/1.96 0.15/2.83 0.21 0.22
Llama-2-13B 0.21/2.07 0.28/3.91 0.19 0.22 0.22/2.19 0.16/2.47 0.22 0.24
Llama-2-70B 0.25/3.07 0.28/3.75 0.22 0.24 0.27/3.55 0.16/2.47 0.24 0.23
GPT-3.5 0.29/4.11 0.36/3.47 0.31 0.23 0.31/3.79 0.31/3.16 0.31 0.29
GPT-4 0.38/4.03 0.39/3.82 0.41 0.29 0.39/4.08 0.32/4.03 0.43 0.35
In-context Learning Llama-2-7B (1-shot) 0.12/2.07 0.19/3.91 0.15 0.16 0.21/2.07 0.16/3.09 0.20 0.19
Llama-2-7B (3-shot) 0.18/2.15 0.21/3.96 0.19 0.18 0.25/2.44 0.18/3.27 0.23 0.20
Llama-2-13B (1-shot) 0.22/2.45 0.26/3.87 0.20 0.19 0.24/3.09 0.15/2.98 0.21 0.23
Llama-2-13B (3-shot) 0.26/3.52 0.31/3.28 0.23 0.24 0.28/2.65 0.21/2.53 0.25 0.26
GPT-3.5 (1-shot) 0.32/4.25 0.38/3.87 0.37 0.28 0.33/3.88 0.30/4.17 0.36 0.34
GPT-3.5 (3-shot) 0.41/4.65 0.42/4.08 0.42 0.31 0.36/4.24 0.36/4.23 0.39 0.36
Vanilla Fine-tuning Llama-2-7B 0.32/4.37 0.38/3.65 0.32 0.27 0.29/3.80 0.30/3.65 0.33 0.31
Llama-2-13B 0.35/4.26 0.41/3.92 0.39 0.29 0.31/4.14 0.29/3.87 0.37 0.35
UrbanKGent Inference Llama-2-7B 0.27/3.05 0.26/4.12 0.28 0.24 0.27/2.87 0.24/3.54 0.26 0.29
Llama-2-13B 0.31/3.87 0.32/3.56 0.35 0.27 0.28/3.24 0.26/3.28 0.31 0.32
Llama-2-70B 0.33/4.28 0.35/4.27 0.33 0.29 0.29/3.80 0.28/4.01 0.32 0.34
GPT-3.5 0.43/4.12 0.46/3.88 0.43 0.34 0.40/4.21 0.39/3.87 0.46 0.41
GPT-4 0.45/4.08 0.48/4.02 0.47 0.42 0.46/4.17 0.41/4.35 0.52 0.43
UrbanKGent-7B 0.46/4.12 0.49/3.97 0.48 0.44 0.49/4.28 0.43/4.58 0.54 0.45
$\uparrow 2.22\%$ $\uparrow 2.08\%$ $\uparrow 2.08\%$ $\uparrow 4.76\%$ $\uparrow 6.52\%$ $\uparrow 4.88\%$ $\uparrow 3.84\%$ $\uparrow 4.66\%$
UrbanKGent-13B 0.52/4.38 0.56/4.13 0.54 0.47 0.53/4.15 0.48/4.42 0.59 0.49
$\uparrow 15.56\%$ $\uparrow 14.29\%$ $\uparrow 14.89\%$ $\uparrow 11.90\%$ $\uparrow 15.22\%$ $\uparrow 17.07\%$ $\uparrow 13.46\%$ $\uparrow 13.95\%$
5.1. Experimental Settings
5.1.1. Dataset
In this work, two sequential tasks (i.e., RTE and KGC) of UrbanKGC are within an open-world setting (i.e., no predefined ontology) (Lu et al., 2023; Ye et al., 2022). Therefore, for the RTE task, every data record is an urban text without the triplet label. For the KGC task, every data record is a quadruple (i.e., head entity name, head entity geometry, tail entity name, tail entity geometry) without the geospatial relation label. We construct the RTE and KGC datasets of NYC and CHI by sampling uniformly from five raw data in Table 1, respectively. As shown in Table 2, we first construct two small datasets (i.e., NYC-Instruct and CHI-Instruct) for instruction fine-tuning and two middle datasets (i.e., NYC and CHI) to validate the performance of the constructed UrbanKGC agent. The remaining data works as the large-scale UrbanKGC dataset (i.e., NYC-Large and CHI-Large) in real-world scenarios shown in Table 7. The three types of datasets are non-overlapping to prevent data leakage. More dataset construction details are in Appendix A.
For each UrbanKGC dataset, we follow Section 4.5 to prompt constructed UrbanKGent to complete RTE and KGC tasks.
5.1.2. Baseline Methods
We provide a comprehensive comparison of our method with existing paradigms (e.g., zero-shot reasoning (Wang et al., 2023a), in-context learning (Wei et al., 2022) and vanilla finetuning (Yao et al., 2023)) on the UrbanKGC tasks.
Pretrained Language Model Baselines. For the RTE task, we utilize RelationPrompt (Chia et al., 2022), an end-to-end generation model for zero-shot RTE. For the KGC task, we fine-tune KG-BERT (Yao et al., 2019) and KG-T5 (Saxena et al., 2022) with the QA pairs constructed from the self-instruct dataset.
LLMs-based Zero-shot Reasoning Methods. We directly prompt the LLMs with basic task definitions to get the answer without training.
LLMs-based In-context Learning Methods. We sample a few-shot QA pairs as demonstrations from the self-instruct dataset as examples and get the answers from the LLMs without training.
Vanilla Fine-tuning Methods. We directly fine-tune the LLMs using the QA pairs constructed from the self-instruct dataset, and then prompt the LLMs with basic task definition without demonstrations.
UrbanKGent Inference Methods. We directly prompt the LLMs using the UrbanKGgent inference pipeline in Section 4.5.
The prompt templates of the above baseline methods are in Appendix B.
5.1.3. Implementation and Detail Settings
In our experiment, we select Llama-2, GPT-3.5 and GPT-4 as the backbone LLM $\mathcal{M}$. We follow the open-source Llama document [^8] https://ai.meta.com/llama/ to deploy Llama-2-7B, Llama-2-13B, and Llama-2-70B in our device. We implement two fine-tuning agent versions (e.g., Llama-2-7B and Llama-2-13B), and all experiments are conducted on eight NVIDIA A800 GPUs. For the GPT-3.5 and GPT-4, we adopt the gpt-3.5-turbo-16k-0613 API and gpt-4-0613 API from the GPT service [^9] https://gpt.hkust-gz.edu.cn/ provided by HKUST (GZ).
5.1.4. Evaluation Protocol
Since the UrbanKGC tasks in this work follow an open-world setting where labels are not visible, the classical metric (e.g., F1 and Hits@10) is not applicable. In this work, we regard evaluation as the binary classification, i.e., if the extracted triplet in RTE task is correct and if the completed relation in KGC task is correct. We follow recent LLMs-based KGC works (Yao et al., 2023) to employ accuracy as an evaluation metric. To make a comprehensive evaluation of the experimental results, we employ the evaluation from two aspects:
Human Evaluation. We employ human annotators to evaluate the results on 200 random samples. For the relational triplet extraction task, we first manually annotate the triplet label for each sample. Then, we manually evaluate the correctness of each triplet (Lu et al., 2023) based on annotation and calculate the accuracy value. For the knowledge graph completion task, we follow (Yao et al., 2023) to manually label the response as correct or wrong, and calculate the accuracy. The detailed human evaluation process is in Appendix D.
GPT Evaluation. Recently, many studies (Zheng et al., 2023; Wang et al., 2023a) adopt LLM-based evaluation for open-domain tasks and empirically demonstrate that GPT-4’s evaluation and human evaluation can be consistent (Cheng et al., 2023). In this work, we also use GPT-4 to evaluate the model performance on the full data to escape intensive labor. Specifically, given an UrbanKGC instruction and results, we prompt GPT-4 to return the confidence score and the justification (i.e., True/False), which will be further used to calculate the accuracy. In this work, the GPT-4’s evaluation has also been demonstrated to be consistent with the human evaluation. The detail is in Appendix D.
Table 4. Effect of different blocks.
Models GPT (acc/confidence) Human (acc)
RTE KGC RTE KGC
${}^{\spadesuit}$ 0.38/4.17 0.42/3.98 0.37 0.34
${}^{\star}$ 0.34/4.06 0.45/4.02 0.34 0.39
${}^{\ddagger}$ 0.45/4.32 0.40/3.97 0.45 0.23
${}^{\dagger}$ 0.44/4.10 0.47/3.85 0.46 0.43
5.2. Main Result (RQ1)
The performance results are reported in Table 3. As can be seen, the constructed agent outperforms all twenty-one baseline models on two UrbanKGC datasets. Specifically, the UrbanKGent-13B achieves (15.56%, 14.29%, 14.89%, and 11.90%) improvements compared with the state-of-the-art GPT-4 with the same inference pipeline on NYC. The improvements on CHI are (15.22%, 17.07%, 13.46%, and 13.95%), respectively. Moreover, the UrbanKGent-7B also achieves comparable performance compared with the GPT-4. Such improvement demonstrates the superiority of the constructed UrbanKGC agent.
Meanwhile, we observe that the zero-shot LLMs perform poorly in the UrbanKGC tasks, even using GPT-4. In addition, although the demonstrations provided by In-context-learning can incorporate the UrbanKGC task information, the performance gain is limited and even leads to performance degradation in smaller LLMs (e.g., Llama-2-7/13B). Besides, we find that fine-tuning LLMs can make obvious improvements in the overall performance. Through vanilla fine-tuning, the Llama-2-7B and Llama-2-13B could achieve comparable performance with GPT-3.5 under the zero-shot reasoning and In-context learning settings.
Moreover, although the LLMs (i.e., Llama-2-7B and Llama-2-13B) using the UrbanKGent inference pipeline perform slightly worse than the vanilla fine-tuning method, they could obtain better performance compared with zero-shot reasoning and In-context learning paradigms. Such results demonstrate the benefit of knowledgeable instruction design and external tool innovation, but also indicate its performance bottleneck (i.e., well-designed UrbanKGent prompting pipeline unable to beat vanilla fine-tuning). As a deeper exploration, our work fills this gap through hybrid instruction fine-tuning, and the fine-tuned UrbanKGC agents, whether 7B or 13B, can achieve state-of-the-art performance in UrbanKGC tasks.
Table 5. Comparison among LLM-based UrbanKGC methods in four ways.
Method
Extra
Knowledge
Require
Fine-tuning
Tool
Invokation
Self
Refinement
ZSL
$\times$ $\times$ $\times$ $\times$
ICL
$\surd$ $\times$ $\times$ $\times$
VFT
$\surd$ $\surd$ $\times$ $\times$
UrbanKGent
Inference
$\surd$ $\times$ $\surd$ $\surd$
UrbanKGent $\surd$ $\surd$ $\surd$ $\surd$
5.3. In-Depth Analysis (RQ2)
We conduct an in-depth analysis of the proposed instruction generation and tool-augmented iterative trajectory refinement module on the NYC dataset. Specifically, for the RTE and KGC task, we validate the effectiveness of each block by comparing the following variants: (1) ${}^{\spadesuit}$ removes knowledgeable instruction template in RTE and KGC task; (2) ${}^{\star}$ removes multi-view design in RTE task; (3) ${}^{\ddagger}$ removes tool invocation; (4) ${}^{\dagger}$ removes iterative trajectory self-refinement. We summarize the results in Table 4, and obtain the following observations.
First, knowledgeable instruction generation contributes to the overall performance of both RTE and KGC tasks. We observe a performance degradation by removing the knowledgeable instruction template. Second, the multi-view instruction design provides the most performance gain, which matches our intuition that the UrbanKG text contains heterogeneous relationships that can be effectively extracted by multi-view prompting design. Third, the tool invocation is very important for the KGC task, as we can observe significant performance degradation after removing the tool invocation. In addition, the iterative trajectory self-refinement brings consistent performance gain for both the RTE and KGC tasks.
Figure 5. The model latency and cost of constructed UrbanKGent-13B and GPT-4 in UrbanKGC tasks. We report the total inference time and cost of 1,000 RTE and KGC tasks.
5.4. Complexity and Effiency Analysis
We make a comparison with the four paradigms to demonstrate the advantages of the constructed agent, which is shown in Table 5. Compared with Zero-shot reasoning (ZSL), In-context Learning (ICL), Vanilla Fine-tuning (VFT), and UrbanKGent Inference, UrbanKGent can incorporate extra urban knowledge, invoke external tools and iteratively self-refine to help better complete UrbanKGC tasks.
Moreover, we also provide efficiency analysis in Figure 5. As can be seen, UrbanKGent-13B achieve lower inference speed in latency and reduce the cost by roughly 20 times in both RTE and KGC tasks. The detail is in Appendix H.
5.5. Agent Application and Deployment (RQ3)
We have deployed a prototype system [^10] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html to facilitate real-world UrbanKGC construction. The service is equipped with the UrbanKGent-13B version and is currently optimized for New York City and Chicago.
As shown in Table 6, compared to the previous UrbanKG benchmark (Ning et al., 2023), we only use roughly one-fifth of the data for constructing the UrbanKGs with the same scale of triplets and entities, and even expanding the variety of relationships to a thousand times the original types. Please refer to Appendix F for more details.
Table 6. Statistics comparision between constructed two UrbanKGs in NYC and CHI and existing UrbanKG dataset and benchmark UUKG (Ning et al., 2023). For relation and data volume, we report the increase and decrease percentages in terms of absolute percentage returns.
Dataset # Entity # Relation # Triplet Data Volume
NYC-Large 265,938 24,052 ($\uparrow 1850\%$) 975,102 39,600 ($\downarrow 597\%$)
CHI-Large 191,473 21,089 ($\uparrow 1622\%$) 610,258 28,378 ($\downarrow 495\%$)
NYC-UUKG 236,287 13 930,240 236,277
CHI-UUKG 140,602 13 564,400 140,577
6. Related work
Our work is related to domain-oriented agent construction, LLMs for knowledge graph construction and urban knowledge graph. We briefly discuss them in this section
6.1. Domain-Oriented Agent Construction
The concept of language agent (Chen et al., 2023) has become very popular recently, and a variety of LLM agents targeting different domains have been proposed. For example, Voyager (Wang et al., 2023c) is constructed for automated game exploration, WebGPT (Gur et al., 2023) is an HTML agent for diverse document understanding tasks, LLMLight (Lai et al., 2023) constructs a language agent for transportation domain, K2 (Deng et al., 2023), GeoGalactica (Lin et al., 2023) and GeoLLM (Li et al., 2023c) propose to re-train language agent for geospatial semantic understanding. In addition, many recent works like Auto-GPT (aut, 2023) and CAMEL (Li et al., 2023a) aim at proposing an autonomous agent framework for agent construction. Nevertheless, there is still no UrbanKGC agent construction framework for the urban computing domain.
6.2. LLMs for Knowledge Graph Construction
Knowledge Graph Construction aims to extract structural information from unstructured text (Ye et al., 2022). Recently, the advent of Large Language Models (LLMs) invigorated the field of Natural Language Processing (NLP). Many studies have begun to explore the potential of LLMs in the domain of knowledge graph construction. For example, (Wei et al., 2023; Li et al., 2023b) finds that transforming the NER and RE task into a multi-turn question-answering dialog could improve the performance of LLMs on the KG construction task. (Xie et al., 2023) explicitly derive syntactic knowledge to guide LLMs to think, which could develop the performance of NER. Despite these LLM-driven knowledge graph construction methods (Yuan et al., 2023; Lu et al., 2023) in general domains being widely investigated, knowledge graph construction in specific domains still remains an open challenge (Zeng and Wang, 2023). There is still no promising knowledge graph construction method in the urban computing domain, and this hinders its potential achievement.
6.3. Urban Knowledge Graph
Urban knowledge graph has been proven useful in various urban tasks, such as traffic flow prediction (Yang and Qi, 2022; Tan et al., 2021; Liu et al., 2021; Zhao et al., 2020), mobility prediction (Wang et al., 2021), site selection (Liu et al., 2022), city profiling (Zhou et al., 2023), crime prediction and so on (Ning et al., 2023). Their common approach involves manually extracting urban entities and defining urban relations to construct an urban knowledge graph. For example, (Wang et al., 2021) construct a dedicated spatiotemporal knowledge graph regarding trajectory and timestamp as entities to improve trajectory prediction and (Liu et al., 2022) construct user check-in relations to help mobility prediction. Nevertheless, existing UrbanKGs heavily rely on manual design, leading to high labor costs and constrained by a limited variety of urban entities and relation categories.
7. Conclusion
In this work, we proposed UrbanKGent, the first automatic UrbanKG construction agent framework with Large Language Models (LLMs). We first constructed a knowledgeable instruction set to adopt LLMs for different UrbanKGC tasks. Then, we proposed a tool-augmented iterative trajectory refinement module to facilitate the instruction tuning of various large language models. Extensive experimental results demonstrate the advancement of UrbanKGent in improving UrbanKGC tasks. We release the UrbanKGent-13B, an agent based on Llama-13B, with lower latency and cost compared with GPT-4 for UrbanKG construction. We hope the open-source UrbanKGent can foster future urban knowledge graph research and broader smart city applications.
8. Limitation and Future Work
This work has limitation on the further application demonstration of construction UrbanKGs, although proposed UrbanKGent-13B could construct a UrbanKG with a thousandfold relationship using only one-fifth of data. In addition, the evaluation method in this work is cost-intensive although GPT evaluation and Human evaluation has been experimentally demonstrated to be consistent. Despite the above limitations, we hope the opensource UrbanKGC agent can foster more extensive UrbanKG research and broad smart city application. In the future, we will derive extra image-modality data to further enrich UrbanKGC.
Appendix A UrbanKGC Construction
Table 7. The detailed statistic of RTE datasets. We report the maximum length, minimum length, and average length of urban text in the RTE dataset.
Dataset Max Length Min Length Avg Length # Records
NYC-Instruct 1,747 68 437 232
CHI-Instruct 1,120 25 408 122
NYC 2,708 51 433 2,089
CHI 1,883 32 445 1,102
NYC-Large 4,277 28 1,167 39,600
CHI-Large 4,306 36 804 28,378
This section presents detailed UrbanKGC dataset statistic information for relational triplet extraction (RTE) and knowledge graph completion (KGC) tasks.
To facilitate the understanding of the constructed RTE dataset, we summarize the distribution of entity types in the four RTE datasets, and the text statistics of entity textual descriptions. As shown in Table 7, the entity distribution and textual statistics of the instruct dataset and test data are similar. In addition, the record in dataset (i.e., NYC-Instruct and CHI-Instruct) used for instruction tuning is not overlapping with that in test dataset (i.e., NYC and CHI). This can avoid potential data leakage issues.
As for the KGC dataset, we provide illustrative visualization of the five RCC relationships (Mani et al., 2010) in Figure 6 for better understanding. Specifically, the disconnected (DC), externally connected (EC), partially overlapping (PO), equal (EQ), tangential and non-tangential proper parts (IN) together depict the basic geospatial relationship between urban entities. Moreover, to facilitate the understanding of the constructed KGC dataset, we visualize the geometry range of the head entity and the tail entity in four KGC datasets. As can be seen in Figure 7, the entity distribution in the instruct dataset (i.e., NYC-Instruct and CHI-Instruct) and test dataset (i.e., NYC and CHI) are similar. Due to the KGC task follows a zero-shot setting, we are unable to provide accurate distribution proportions of the 5 RCC relationships in the four datasets. However, it is intuitive can be seen that all four datasets contain the five RCC relationships shown in Figure 6, which provides a guarantee for the practical significance of KGC task in this work.
Figure 6. Given two geo-entities X and Y, the illustrative visualization of five types of RCC relationships. In this work, we consider entities with coordinate geometry to a circles with very small radii.
Appendix B Instruction Template
B.1. Instruction Template in UrbanKGent
This section presents the detailed instruction template of the proposed UrbanKGent framework. Specifically, Figure 8(a-b) provides the detailed instruction template of knowledgeable instruction generation module and tool-based trajectory augmentation. The iterative trajectory self-refinement module is achieved by trajectory verifier and updater in Figure 8(c).
Figure 7. The geometry range visualization of the head entity and tail entity of four KGC datasets. The horizontal and vertical coordinates are longitude and latitude, respectively. The blue and red polygons stand for entities with the polygon geometry, the purple line string stands for the entities with linestring geometry and the green point is for the coordinate entities.
B.2. Instruction Template in Baselines
We provide the detailed instruction template of all the baseline models in this work.
LLMs-based Zero-shot Reasoning Methods. We only provide task descriptions for the zero-shot reasoning method. The detailed instruction template is shown in Figure 9.
LLMs-based In-context Learning Methods. We first construct several few-shot demonstrations via the chain-of-thought prompting techniques (Wei et al., 2022), which is popular for automatic demonstration generation. The detailed instruction template is shown in Figure 9. Then, we add these demonstrations before the test question, and the detailed instruction template is shown in Figure 9.
Vanilla Fine-tuning Methods. We only provide task descriptions for vanilla fine-tuning methods without demonstrations. The detailed instruction template is shown in Figure 9.
UrbanKGent Inference Methods. These baseline models follow the same inference pipeline shown in Figure 8.
Figure 8. The reasoning process of UrbanKGent-13B and the detailed instruction template for UrbanKent inference pipeline. The content in the gray box is the instruction, and the content in the green box is the agent’s response.
Figure 9. The instruction template of zero-shot reasoning, In-context learning and vanilla fine-tuning baselines.
Table 8. The detailed geospatial tool name, description, interface input, and output. Each tool interface is implemented by python and it is self-programmed by GPT-4.
Tool name Tool description Input Output
Geohash Geohash encoding Geometry Geohash code (8-bit)
Distance Calculate the distance between two geo entities. Geometry 1, Geometry 2 Distance value (km)
Point2Polygon Identify if a point belongs to a polygon Point geometry, polygon geometry True/False
Point4Linestring Identify if a point intersects a linestring Point geometry, linestring geometry True/False
Linestring2Polygon Identify if a linestring belongs to a polygon Linestring geometry, polygon geometry True/False
Linestring4Polygon Identify if a linestring intersects a polygon Linestring geometry, polygon geometry True/False
Polygon2Polygon Identify if a polygon belongs to a polygon Polygon geometry, polygon geometry True/False
Polygon4Polygon Identify if a polygon intersects a polygon Polygon geometry, polygon geometry True/False
Appendix C Geospatial Toolkit
Table 9. Illustrative RTE evaluation example when we utilize human evaluation and GPT evaluation method. We calculate the accuracy by counting the proportion of true triplets. The label for GPT evaluation method is invisible.
Type Urban text Results Label
Number of the
true triplet
Number of the
false triplet
Human
Evaluation
Columbia University is a private Ivy league research university in New York City. <Columbia University, locate-in, New York City >
<Columbia University,
locate-in, New York City >
1 0
GPT
Evaluation
- 1 0
This work constructs a geospatial reasoning toolkit (eight interfaces in total in Table 8) by asking GPT-4 for self-programming. We obtained geospatial toolkit for point, linestring, and polygon geometry objects, supporting distance calculation and Geohash encoding for two geographical objects, and can also determine basic spatial relationships between two geometric objects, such as containment and intersection. For each interface, we provide its name and function description (e.g., geohash encoding in Table 8 and Figure 8b) into the instruction template.
Appendix D Evaluation
This section presents the detailed evaluation process and examples to help the reader have a better understanding. Moreover, we provide a comprehensive analysis to demonstrate why GPT and Human evaluations are highly aligned.
D.1. Human Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 9. Given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, the human annotators are required to first label the triplet contained in this urban text, described as triplet ¡Columbia University, locate-in, New York City¿. Based on the label, then, they are instructed to evaluate how many true triplets and false triplets in the results from the models. Finally, they will fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). We will calculate the accuracy of results based on these annotated forms.
For the knowledge graph completion (KGC) task, given the head entity ¡Columbia University, Point (40.8075, 73.9626)¿ and the tail entity ¡Empire State Building, Point (40.7484, 73.9857)¿, the human annotators are first required to complete their missing geospatial relationship from the five relation candidate (i.e., DC, EC, PO, EQ and IN). Specifically, the annotation could be achieved by manually visualizing the location of two entities given on the map and following the RCC relation rule in Figure 6 to determine their geographical relationships. Finally, they will fill out the evaluation results (i.e., True/False). We will calculate the accuracy based on these evaluation results.
D.2. GPT Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 9. Given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, and the model extraction results (i.e., ¡Columbia University, locate-in, New York City¿), we directly prompt GPT-4 to fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). Then, the accuracy could be obtained based on these self-evaluated results.
For the knowledge graph completion (KGC) task, given the head entity ¡Columbia University, Point (40.8075, 73.9626)¿ and the tail entity ¡Empire State Building, Point (40.7484, 73.9857)¿ and completed geospatial relationship, we directly prompt GPT-4 the justify if the results are correct. Specifically, we will first explicitly call eight external geospatial tools in Table 8, and combine all these eight calculation results into the prompt. Then we feed these prompts into GPT-4 to help it make a comparable accurate evaluation. Finally, we obtain the evaluation results (i.e., True/False) for every record. We will calculate the accuracy based on these evaluation results.
However, such an evaluation process carries risks as we cannot guarantee the infallibility of GPT-4. To ensure the validity of the evaluation, we have conducted extensive experiments to demonstrate that the GPT evaluation method aligns well with human evaluation.
D.3. Evaluation Consistency
We conduct experiments to assess the consistency between GPT-4’s evaluation results and human evaluation results. In this work, we follow GPTScore (Fu et al., 2023a) to use Spearman coefficient (Zar, 2005) to investigate the correlation between GPT evaluation and human evaluation. Based on the evaluation results in Tabel 3, we can conduct correlation analysis separately by the types of model and paradigm.
Comparison of LLM Backbones. As shown in Figure 10, we observe that the Spearman correlation coefficients on all LLM backbones are greater than $0.8$, which means the human evaluation and GPT evaluation on these five LLM backbones is highly correlated.
Comparison of LLM Paradigm. Moreover, we also conduct correlation under different LLM paradigms for UrbanKGC tasks. As can be seen in Figure 11, all four types of paradigm have a higher Spearman value above 0.85. Therefore, deriving the GPT evaluation method is also applicable to different LLM paradigms when finishing UrbanKGC tasks as its evaluation aligns well with that of the human being.
Analysis of Evaluation Repeat. Nevertheless, the above correlation analysis is based on results in Table 3, which are obtained after only single evaluation. To deeply understand the evaluation mechanism, We repeatedly instruct GPT-4 to generate 1, 3, 5, and 10 evaluation results to understand the variance of the evaluation method. To save the cost, we perform the above analysis using UrbanKGent-7B only on the NYC dataset. The results are shown in Table 10, although the coefficient value will decrease as the number of repeated experiments increases. However its value has always remained within a high correlation range (roughly 0.85).
Table 10. The average Spearman correlation value between human evaluations and GPT-4 evaluations. ”Repeat X times” refers to instructing GPT-4 to generate judgments X times, and adopting the answer that appears most frequently (e.g., True/False for the KGC task and Number of the true triplet for the RTE task) as the final decision.
Task Repeat Once Repeat 3 times Repeat 5 times Repeat 10 times
RTE 98.56% 87.31% 86.42% 85.27%
KGC 94.86% 91.23% 90.09% 91.24%
Appendix E Case Study
As shown in Figure 8, we present the detailed reasoning process of constructed UrbanKGent-13B when finishing urban relational triplet extraction and knowledge graph completion task. Since the iterative self-refinement process contains excessive text, we display the reasoning process of only one iteration.
Figure 10. The Spearman correlation between the GPT evaluation and human’s evaluation under five different LLM backbones (i.e., Llama-2-7B, Llama-2-13B, Llama-2-70B, GPT-3.5 and GPT-4).
Figure 11. The Spearman correlation between the GPT evaluation and human’s evaluation under four LLM paradigms (i.e., Zero-shot learning (ZSL), In-context learning (ICL), Vanilla fine-tuning (VFL) and UrbanKGent inference).
Appendix F Prototype System
We have deployed the constructed UrbanKGent-13B as a prototype system to offer urban knowledge graph construction service. The project service could be found in Github [^11] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html. Figure 12 shows the screenshot of our demo system. Specifically, for the user, the system provides UrbanKGC service including relational triplet extraction and knowledge graph completion task. Following the instruction template proposed in this paper, the user can input urban text in the web shown in Figure 12 and the web will return extracted triplet. The user can also input two entities name and their geometry value, and the web serive will return completed geospatial relationship between them.
Appendix G Evaluation on Larger Dataset
As shown in Table 6, we derive UrbanKGent-13B for urban knowledge graph construction using constructed large-scale dataset in NYC and CHI. Specifically, we directly use the text record in NYC-Large and CHI-Large for the relational triplet extraction task. Then, we randomly sample the head-tail entity pairs (both of head and tail entities contain geometry information) from these triplets for knowledge graph completion. Nevertheless, iterating all head-tail entity pairs is a time-consuming task, so we just construct a KGC dataset consistent with the sacle of the RTE dataset. By performing RTE and KGC tasks, we obtain two large-scale UrbanKGs shown in Table 6. Compared with existing dataset and benchmark UUKG (Ning et al., 2023), we can clearly observe that our agent can only use one-fifth data to construct the UrbanKG with the same scale entities and triplets, but extend the relationship types to a thousand times. It is worth noting that all construction process is completed by a LLM agent without any mannul effort. We think it is the core of this work.
Figure 12. The prototype system of UrbanKGent.
Moreover, we also report the performance of the RTE and KGC in NYC-Large dataset. To save the cost of GPT self-evaluation, we only choose the best approaches in ZSL, ICL, VFT and UrbanKGent Inference. The experimental results are shown in Table 11. As can be seen, the fine-tuned UrbanKGC agent, whether 7B or 13B version, could achieve state-of-art performance on UrbanKGC tasks.
Table 11. The experimental results of relational triplet extraction (RTE) and knowledge graph completion (KGC) on the NYC-Large dataset. To save the cost, We choose the best approaches in zero-shot-learning (ZSL), In-context-learning (ICL), Vanilla Fine-tuning (VFT), and UrbanKGent Inference setting.
Models GPT(acc/confidence) Human(acc)
RTE KGC RTE KGC
ZSL 0.36/4.01 0.45/3.95 0.42 0. 31
ICL 0.39/4.36 0.48/4.15 0.48 0.39
VFT 0.37/4.05 0.46/3.98 0.45 0.35
UrbanKGent
Inference
0.43/4.13 0.51/3.87 0.51 0.43
UrbanKGent-7B 0.44/4.27 0.50/4.07 0.53 0.44
UrbanKGent-13B 0.46/4.56 0.52/3.67 0.55 0.46
Appendix H Complexity and Effiency Analysis
We make a comparison with the four paradigms to demonstrate the advantages of the constructed agent, which is shown in Table 5. Compared with Zero-shot reasoning (ZSL), In-context Learning (ICL), Vanilla Fine-tuning (VFT), and UrbanKGent Inference, UrbanKGent can incorporate extra urban knowledge, invoke external tools and iteratively self-refine to help better complete UrbanKGC tasks.
Moreover, we also provide comprehensive efficiency analysis to show the latency and cost of different models when completing UrbanKGC tasks. Specifically, we report the total inference time and cost [^12] We subscribe to the NVIDIA A800 computing resources and GPT service from HKUST(GZ). Following the standard price instruction of HKUST(GZ), we could calculate cost of GPT-based baselines or Llama-based baselines. But in order to protect privacy of the university, we do not provide detailed prices information here. of each method completing with 1,000 RTE and KGC tasks. For the cost of GPT-4 service, we first count the number of prompt token and completion token spent on 1,000 RTE and KGC tasks, and then calculate the cost based on billing standards. As for the cost of UrbanKGent-13B, we first count the GPU running time spent on 1,000 RTE and KGC tasks, and then calculate the cost based on the A800 charging standard provided by HKUST(GZ).